The Jones Company has just completed the third year of a​ five-year MACRS recovery period for a piece of equipment it originally purchased for $295,000.

a. What is the book value of the​ equipment?

b. If Jones sells the equipment today for $185,000 and its tax rate is 35%​, what is the​ after-tax cash flow from selling​ it?

c. Just before it is about to sell the​ equipment, Jones receives a new order. It can take the new order if it keeps the old equipment. Is there a cost to taking the order and if​ so, what is​ it? Explain.​ (Assume the new order will consume the remainder of the​ machine's useful​ life.)

Note​:

Assume that the equipment is put into use in year 1.
Answer #1

1: Book value of the​ equipment= Initial cost -Depreciation till date

= 295000- 295000*(20%+32%+19.2%)

= 295000-210040

= 84960

2: Gain from sale= 185000-84960 = 100040

After tax cash flow= Selling price-Tax

=185000-35%*100040

= 149986

3: The cost is the opportunity cost of selling price lost due to keeping the old equipment.